OpenStudy (anonymous):

-10 + sqrt x+8= -4 can someone take me step by step on how to solve this ?

OpenStudy (jdoe0001):

\(\bf -10 + \sqrt{x}+8= -4 \qquad\qquad {\color{brown}{ +10\ and -8}} \\ \quad \\ \cancel{ -10 {\color{brown}{ +10}} }+ \sqrt{x}\cancel{ +8{\color{brown}{ -8}} }= -4{\color{brown}{ +10}}{\color{brown}{ -8}}\implies \sqrt{x}=-2\qquad \qquad {\color{brown}{ ^2}} \\ \quad \\ (\sqrt{x})^{\color{brown}{ 2}}=(-2)^{\color{brown}{ 2}}\implies x=+4\)

OpenStudy (anonymous):

the sqrt goes over both the x and the 8

OpenStudy (jdoe0001):

then is the same thing... you 'd just add +10 only then and then square both sides

OpenStudy (jdoe0001):

the idea is just to isolate the root on the left-hand-side then square both sides

OpenStudy (anonymous):

oh okay thanks

OpenStudy (jdoe0001):

yw